# Aperiodic sets of three types of convex polygons

### v1

@@ Appendix D.1 Preparation
The restriction on the number of types of congruent diagrams is imposed because the method of creating the ASP introduced in the study of this appendix would create a variety of $ASPcvx(n)$ with $n\geq 4$ ($ASPcvx(n\geq 4,m\geq 4)$) that do not satisfy this restriction. In addition, considering that only a small number of types of prototype tiles included in an ASP has been explored so far, we thought that it would be meaningful to consider cases that satisfy this restriction.
@@ Appendix D.2 Dividing Tile(1,1)11(1,1)( 1 , 1 ) into convex pentagons and convex hexagons
Tile(1,1)11(1,1)( 1 , 1 ), corresponding to $a=b=1$ in Tile$(a,b)$ shown in [9], can generate a periodic tiling if the use of reflected tiles is allowed during the tiling generation process (see Figure 2(b)). However, as shown in [10], Tile(1,1)11(1,1)( 1 , 1 ) can only generate non-periodic tiling if and only if it does not allow the use of reflected tiles (see Figure 3). Using this property, $ASPcvx(n)$ can be obtained from Tile(1,1)11(1,1)( 1 , 1 ).
@@
Figure. 27: Cases of $\alpha+\beta=180^{\circ}$, $\beta=120^{\circ}$ or $\alpha=165^{\circ}$.
Figure. 28: Cases of $\alpha+\beta=210^{\circ}$, $\alpha=135^{\circ}$ or $\beta=90^{\circ}$.
As shown in Figure 26, by changing the parameters $\alpha$ and $\beta$, the shape of the five convex polygons inside Tile(1,1)11(1,1)( 1 , 1 ) changes as point $P$ moves on the bisector $X_{5}Q$ of the interior angle $X_{5}$ and point $T$ moves on the bisector $X_{10}S$ of the interior angle $X_{10}$ (the edge corresponding to the green line in Figure 26 has its length changed by changing $\alpha$ and $\beta$). Based on the change, the following cases appear inside Tile(1,1)11(1,1)( 1 , 1 ) from a viewpoint that does not distinguish between anterior and posterior sides: “one type of convex pentagon and two types of convex hexagons,” “two types of convex pentagons and one type of convex hexagon,” and “three types of convex pentagons.”
[…]
Let $S_{t}(\alpha,\beta)$ be the tile set of {$CP(X_{14})$, $CP(X_{3})$, $CP(X_{7})$, $CP(X_{9})$, $CP(X_{11})$} for parameters $\alpha$ and $\beta$. The convex polygons contained in $S_{t}(\alpha,\beta)$ become convex pentagonal or hexagonal monotiles if $\alpha$ and $\beta$ satisfy certain conditions, as listed in Table 1 [3, 11, 12, 18]. All cases in in Table 1 contain the convex pentagonal or hexagonal monotiles that can form periodic tilings without using reflected tiles (see Appendix A and Appendix B)111111 The case of $\beta=120^{\circ}$ and $105^{\circ}\leq\alpha<165^{\circ}$ also contains the convex pentagonal monotiles belonging only to the Type 2 family which require the use of reflected tiles during the tiling generation process. However, the tile set in the case contains the anterior and posterior sides of convex pentagonal monotiles belonging to the Type 2 family. .
@@
$\alpha+\beta=210^{\circ}$ Convex hexagonal monotile belonging to the Type 1 family Figure 28(a)
@@
$\alpha=135^{\circ}$ Convex hexagonal monotiles belonging to the Type 3 family Figure 28(b)
@@
$\beta=90^{\circ}$ Convex pentagonal monotiles belonging to both the Type 2 and Type 4 families Figure 28(b)
@@
$\alpha=135^{\circ}$ and $\beta=90^{\circ}$ Convex hexagonal monotile belonging to the Type 3 family Convex hexagonal monotiles belonging to the Type 3 family Convex pentagonal monotiles belonging to both the Type 2 and Type 4 families Figure 28(b)
@@
$\beta=120^{\circ}$ and $105^{\circ}\leq\alpha<165^{\circ}$ Convex pentagonal monotile belonging to the Type 1 family Convex pentagonal monotiles belonging to the Type 2 family Figure 27(b)
@@
Appendix D.3 ASP satisfying the restriction on the number of types of congruent diagrams created from Tile(1,1)11(1,1)( 1 , 1 )
@@
If the convex polygons contained in $S_{t}(\alpha,\beta)$ are not convex pentagonal or hexagonal monotiles, we can create candidates of $ASPcvx(4,3)$ and $ASPcvx(5,3)$ from Tile(1,1)11(1,1)( 1 , 1 ). We refer to it as “candidate” because if we generate tilings with convex polygons contained in $S_{t}(\alpha,\beta)$, we have to make sure that there exist only non-periodic tilings using the shape of Tile(1,1)11(1,1)( 1 , 1 ), if and only if all types of convex polygons contained in $S_{t}(\alpha,\beta)$ are used. Therefore, we will confirm it with the following steps:
@@
  • Step 1. Check that the convex polygons contained in $S_{t}(\alpha,\beta)$ are not monotile.
@@
  • Step 2. If it uses more than one type (but not all) of the convex polygons contained in $S_{t}(\alpha,\beta)$, it cannot admit tilings.
@@
  • Step 3. Check that with all the convex polygons contained in $S_{t}(\alpha,\beta)$, they cannot admit tilings in any combination other than the shape of Tile(1,1)11(1,1)( 1 , 1 ), as shown in Figure 26.
@@
For example, we confirmed that $S_{t}(\alpha=105^{\circ},\beta=75^{\circ})$, which includes five convex pentagons created by dividing Tile(1,1)11(1,1)( 1 , 1 ) shown in Figure 29, satisfies Steps 1–3. As shown in Figure 29, when $\alpha=105^{\circ}$ and $\beta=75^{\circ}$ are used, $CP(X_{3})$ and $CP(X_{7})$ are convex pentagons with line symmetry and cannot distinguish between the anterior and posterior sides. Therefore, $S_{t}(\alpha=105^{\circ},\beta=75^{\circ})$ is $ASPcvx(4,3)$. In $S_{t}(\alpha,\beta)$, we conjecture that Figure 29 is the only case where the ASP comprises of four types of convex pentagons, and the ASP comprises of three types of convex pentagons from a viewpoint that does not distinguish between the anterior and posterior sides.
@@
In Steps 1–3, we briefly describe some of the key points for identifying cases where the convex polygons in Figure 29 cannot form tilings. We focused on the interior angles of the convex polygons at $75^{\circ}$, $105^{\circ}$, $135^{\circ}$, and $165^{\circ}$. To form tilings, two or more tile vertices must be concentrated at one point, and the sum of the interior angles of the vertices at the concentration point must be $180^{\circ}$ or $360^{\circ}$. Therefore, as there are no convex pentagons with interior angles of $60^{\circ}$ or less, $135^{\circ}$ and $165^{\circ}$ cannot be used in a concentration with a sum of $180^{\circ}$. We also observe that an even number of vertices with interior angles of $75^{\circ}$, $105^{\circ}$, $135^{\circ}$, and $165^{\circ}$ must always be concentrated at one point because the place value of one is “five.” Based on these considerations, it was concluded that the concentration of vertices with an interior angle of $135^{\circ}$ was impossible, except for the arrangements using $CP(X_{9})$ and $CP(X_{11})$, as shown in Figure 30(a). Subsequently, note the point in Figure 30(a) where the two vertices have an internal angle of $135^{\circ}$. If a tiling is formed, the only method to obtain a sum of $360^{\circ}$ is to place the vertex of the tile with an interior angle of $90^{\circ}$ at the concentrating point. Figure 30(b) shows the only combination that can form a structure in which the tiling does not collapse. Then, for the positions with vertices having internal angles of $75^{\circ}$ and $165^{\circ}$ circled in blue in the cluster in Figure 30(b), the only combination that can construct a structure in which the tiling does not collapse using $S_{t}(\alpha=105^{\circ},\beta=75^{\circ})$ is that of dividing Tile(1,1)11(1,1)( 1 , 1 ) (using the shape of Tile(1,1)11(1,1)( 1 , 1 )), as shown in Figure 29.
@@
Other candidates for $ASPcvx(4,3)$ and $ASPcvx(5,3)$ related to $S_{t}(\alpha,\beta)$ are as follows:
@@
      Case 1. When $\alpha=105^{\circ}$, $75^{\circ}<\beta<120^{\circ}$, $\beta\neq 90^{\circ}$, and $\alpha+\beta\neq 210^{\circ}$ (Figures 26 and 29), there is a “set of three types of convex pentagons ($CP(X_{3})$, $CP(X_{7})$, $CP(X_{9})$, $CP(X_{11})$) and one type of convex hexagon ($CP(X_{14})$),” and, from the viewpoint that does not distinguish between anterior and posterior sides, it is a “set of two types of convex pentagons and one type of convex hexagon.”
@@
      Case 2. When $\alpha+\beta=180^{\circ}$ and $\alpha\neq 105^{\circ}$ (Figure 27(a)), there is a “set of three types of convex pentagons ($CP(X_{14})$, $CP(X_{9})$, $CP(X_{11})$) and two types of convex hexagons ($CP(X_{3})$, $CP(X_{7})$),” and, from the viewpoint that does not distinguish between anterior and posterior sides, it is a “set of two types of convex pentagons and one type of convex hexagon.”
@@
      Case 3. When $\alpha=165^{\circ}$, $65.104^{\circ}<\beta<120^{\circ}$, and $\beta\neq 90^{\circ}$ (Figure 27(c)), there is a “set of three types of convex pentagons ($CP(X_{14})$, $CP(X_{9})$, $CP(X_{11})$) and two types of convex hexagons ($CP(X_{3})$, $CP(X_{7})$),” and, from the viewpoint that does not distinguish between anterior and posterior sides, it is “set of two types of convex pentagons and one type of convex hexagon.”
@@
      Case 4. When $105^{\circ}<\alpha<165^{\circ}$, $\alpha\neq 135^{\circ}$, $65.104^{\circ}<\beta<120^{\circ}$, $\beta\neq 90^{\circ}$, $\alpha+\beta>180^{\circ}$, and $\alpha+\beta\neq 210^{\circ}$ (Figure 26), there is a “set of two types of convex pentagons ($CP(X_{9})$, $CP(X_{11})$) and three types of convex hexagons ($CP(X_{14})$, $CP(X_{3})$, $CP(X_{7})$),” and, from the viewpoint that does not distinguish between anterior and posterior sides, it is a “set of one type of convex pentagon and two types of convex hexagon.”
As can be observed from the parameters $\alpha$ and $\beta$, in contrast to the case shown in Figure 29, the shapes of the convex polygons in the tile sets in Cases 1–4 are not uniquely determined. Tilings using the shape of Tile(1,1)11(1,1)( 1 , 1 ) in the tile sets of Cases 1 and 4 were edge-to-edge tilings with convex polygons. Tilings using the shape of Tile(1,1)11(1,1)( 1 , 1 ) in the tile sets of Cases 2 and 3 were non-edge-to-edge tilings with convex polygons.

### v3

@@ Appendix D.1 Preparation
The restriction on the number of types of congruent diagrams is imposed because the method of creating the ASP introduced in the study of this appendix would create a variety of $ASPcvx(n)$ with $n\geq 4$ ($ASPcvx(n\geq 4,m\geq 4)$) that do not satisfy this restriction. In addition, considering that only a small number of types of prototype tiles contained in an ASP has been explored so far, we thought that it would be meaningful to consider cases that satisfy this restriction.
@@ Appendix D.2 Dividing Tile(1,1)11(1,1)( 1 , 1 ) into convex pentagons and convex hexagons
Tile(1,1)11(1,1)( 1 , 1 ), corresponding to $a=b=1$ in Tile$(a,b)$ shown in [9], can generate a periodic tiling if the use of reflected tiles is allowed during the tiling generation process (see Figure 2). However, as shown in [10], Tile(1,1)11(1,1)( 1 , 1 ) can only generate non-periodic tiling if and only if it does not allow the use of reflected tiles (see Figure 3). Using this property, $ASPcvx(n)$ can be obtained from Tile(1,1)11(1,1)( 1 , 1 ).
@@
As shown in Figure 29, by changing the parameters $\delta$ and $\epsilon$, the shape of the five convex polygons inside Tile(1,1)11(1,1)( 1 , 1 ) changes as point $P$ moves on the bisector $X_{5}Q$ of the interior angle $X_{5}$ and point $T$ moves on the bisector $X_{10}S$ of the interior angle $X_{10}$ (the edge corresponding to the green line in Figure 29 has its length changed by changing $\delta$ and $\epsilon$). Based on the change, the following cases appear inside Tile(1,1)11(1,1)( 1 , 1 ) from a viewpoint that does not distinguish between anterior and posterior sides: “one type of convex pentagon and two types of convex hexagons,” “two types of convex pentagons and one type of convex hexagon,” and “three types of convex pentagons.”
[…]
Let $S_{t}(\delta,\epsilon)$ be the tile set of {$CP(X_{14})$, $CP(X_{3})$, $CP(X_{7})$, $CP(X_{9})$, $CP(X_{11})$} for parameters $\delta$ and $\epsilon$. The convex polygons contained in $S_{t}(\delta,\epsilon)$ become convex pentagonal or hexagonal monotiles if $\delta$ and $\epsilon$ satisfy certain conditions, as listed in Table 1 [3, 11, 12, 18]. All cases in in Table 1 contain the convex pentagonal or hexagonal monotiles that can form periodic tilings without using reflected tiles (see Appendix A and Appendix B)121212 The case of $\epsilon=120^{\circ}$ and $105^{\circ}\leq\delta<165^{\circ}$ also contains the convex pentagonal monotiles belonging only to the Type 2 family which require the use of reflected tiles during the tiling generation process. However, the tile set in the case contains the anterior and posterior sides of convex pentagonal monotiles belonging to the Type 2 family. .
@@
Figure. 30: Cases of $\delta+\epsilon=180^{\circ}$, $\epsilon=120^{\circ}$ or $\delta=165^{\circ}$.
Figure. 31: Cases of $\delta+\epsilon=210^{\circ}$, $\delta=135^{\circ}$ or $\epsilon=90^{\circ}$.
@@
$\delta+\epsilon=210^{\circ}$ Convex hexagonal monotile belonging to the Type 1 family Figure 31(a)
@@
$\delta=135^{\circ}$ Convex hexagonal monotiles belonging to the Type 3 family Figure 31(b)
@@
$\epsilon=90^{\circ}$ Convex pentagonal monotiles belonging to both the Type 2 and Type 4 families Figure 31(b)
@@
$\delta=135^{\circ}$ and $\epsilon=90^{\circ}$ Convex hexagonal monotile belonging to the Type 3 family Convex hexagonal monotiles belonging to the Type 3 family Convex pentagonal monotiles belonging to both the Type 2 and Type 4 families Figure 31(b)
@@
$\epsilon=120^{\circ}$ and $105^{\circ}\leq\delta<165^{\circ}$ Convex pentagonal monotile belonging to the Type 1 family Convex pentagonal monotiles belonging to the Type 2 family Figure 30(b)
@@
Appendix D.3 ASPs satisfying the restriction on the number of types of congruent diagrams created from Tile(1,1)11(1,1)( 1 , 1 )
@@
If the convex polygons contained in $S_{t}(\delta,\epsilon)$ are not convex pentagonal or hexagonal monotiles, we can create candidates of $ASPcvx(4,3)$ and $ASPcvx(5,3)$ from Tile(1,1)11(1,1)( 1 , 1 ). We refer to it as “candidate” because if we generate tilings with convex polygons contained in $S_{t}(\delta,\epsilon)$, we have to make sure that there exist only non-periodic tilings using the shape of Tile(1,1)11(1,1)( 1 , 1 ), if and only if all types of convex polygons contained in $S_{t}(\delta,\epsilon)$ are used. Therefore, we will confirm it with the following steps:
@@
  • Step 1. Check that the convex polygons contained in $S_{t}(\delta,\epsilon)$ are not monotile.
@@
  • Step 2. If it uses more than one type (but not all) of the convex polygons contained in $S_{t}(\delta,\epsilon)$, it cannot admit tilings.
@@
  • Step 3. Check that with all the convex polygons contained in $S_{t}(\delta,\epsilon)$, they cannot admit tilings in any combination other than the shape of Tile(1,1)11(1,1)( 1 , 1 ), as shown in Figure 29.
@@
For example, we confirmed that $S_{t}(\delta=105^{\circ},\epsilon=75^{\circ})$, which contains five convex pentagons created by dividing Tile(1,1)11(1,1)( 1 , 1 ) shown in Figure 32, satisfies Steps 1–3. As shown in Figure 32, when $\delta=105^{\circ}$ and $\epsilon=75^{\circ}$ are used, $CP(X_{3})$ and $CP(X_{7})$ are convex pentagons with line symmetry and cannot distinguish between the anterior and posterior sides. Therefore, $S_{t}(\delta=105^{\circ},\epsilon=75^{\circ})$ is $ASPcvx(4,3)$. In $S_{t}(\delta,\epsilon)$, we conjecture that Figure 32 is the only case where the ASP comprises of four types of convex pentagons, and the ASP comprises of three types of convex pentagons from a viewpoint that does not distinguish between the anterior and posterior sides.
@@
In Steps 1–3, we briefly describe some of the key points for identifying cases where the convex polygons in Figure 32 cannot form tilings. We focused on the interior angles of the convex polygons at $75^{\circ}$, $105^{\circ}$, $135^{\circ}$, and $165^{\circ}$. To form tilings, two or more tile vertices must be concentrated at one point, and the sum of the interior angles of the vertices at the concentration point must be $180^{\circ}$ or $360^{\circ}$. Therefore, as there are no convex pentagons with interior angles of $60^{\circ}$ or less, $135^{\circ}$ and $165^{\circ}$ cannot be used in a concentration with a sum of $180^{\circ}$. We also observe that an even number of vertices with interior angles of $75^{\circ}$, $105^{\circ}$, $135^{\circ}$, and $165^{\circ}$ must always be concentrated at one point because the place value of one is “five.” Based on these considerations, it was concluded that the concentration of vertices with an interior angle of $135^{\circ}$ was impossible, except for the arrangements using $CP(X_{9})$ and $CP(X_{11})$, as shown in Figure 33(a). Subsequently, note the point in Figure 33(a) where the two vertices have an internal angle of $135^{\circ}$. If a tiling is formed, the only method to obtain a sum of $360^{\circ}$ is to place the vertex of the tile with an interior angle of $90^{\circ}$ at the concentrating point. Figure 33(b) shows the only combination that can form a structure in which the tiling does not collapse. Then, for the positions with vertices having internal angles of $75^{\circ}$ and $165^{\circ}$ circled in blue in the cluster in Figure 33(b), the only combination that can construct a structure in which the tiling does not collapse using $S_{t}(\delta=105^{\circ},\epsilon=75^{\circ})$ is that of dividing Tile(1,1)11(1,1)( 1 , 1 ) (using the shape of Tile(1,1)11(1,1)( 1 , 1 )), as shown in Figure 32.
@@
Other candidates for $ASPcvx(4,3)$ and $ASPcvx(5,3)$ related to $S_{t}(\delta,\epsilon)$ are as follows:
@@
      Case 1. When $\delta=105^{\circ}$, $75^{\circ}<\epsilon<120^{\circ}$, $\epsilon\neq 90^{\circ}$, and $\delta+\epsilon\neq 210^{\circ}$ (Figures 29 and 32), there is a “set of three types of convex pentagons ($CP(X_{3})$, $CP(X_{7})$, $CP(X_{9})$, $CP(X_{11})$) and one type of convex hexagon ($CP(X_{14})$),” and, from the viewpoint that does not distinguish between anterior and posterior sides, it is a “set of two types of convex pentagons and one type of convex hexagon.”
@@
      Case 2. When $\delta+\epsilon=180^{\circ}$ and $\delta\neq 105^{\circ}$ (Figure 30(a)), there is a “set of three types of convex pentagons ($CP(X_{14})$, $CP(X_{9})$, $CP(X_{11})$) and two types of convex hexagons ($CP(X_{3})$, $CP(X_{7})$),” and, from the viewpoint that does not distinguish between anterior and posterior sides, it is a “set of two types of convex pentagons and one type of convex hexagon.”
@@
      Case 3. When $\delta=165^{\circ}$, $65.104^{\circ}<\epsilon<120^{\circ}$, and $\epsilon\neq 90^{\circ}$ (Figure 30(c)), there is a “set of three types of convex pentagons ($CP(X_{14})$, $CP(X_{9})$, $CP(X_{11})$) and two types of convex hexagons ($CP(X_{3})$, $CP(X_{7})$),” and, from the viewpoint that does not distinguish between anterior and posterior sides, it is “set of two types of convex pentagons and one type of convex hexagon.”
@@
      Case 4. When $105^{\circ}<\delta<165^{\circ}$, $\delta\neq 135^{\circ}$, $65.104^{\circ}<\epsilon<120^{\circ}$, $\epsilon\neq 90^{\circ}$, $\delta+\epsilon>180^{\circ}$, and $\delta+\epsilon\neq 210^{\circ}$ (Figure 29), there is a “set of two types of convex pentagons ($CP(X_{9})$, $CP(X_{11})$) and three types of convex hexagons ($CP(X_{14})$, $CP(X_{3})$, $CP(X_{7})$),” and, from the viewpoint that does not distinguish between anterior and posterior sides, it is a “set of one type of convex pentagon and two types of convex hexagon.”
As can be observed from the parameters $\delta$ and $\epsilon$, in contrast to the case shown in Figure 32, the shapes of the convex polygons in the tile sets in Cases 1–4 are not uniquely determined. Tilings using the shape of Tile(1,1)11(1,1)( 1 , 1 ) in the tile sets of Cases 1 and 4 were edge-to-edge tilings with convex polygons. Tilings using the shape of Tile(1,1)11(1,1)( 1 , 1 ) in the tile sets of Cases 2 and 3 were non-edge-to-edge tilings with convex polygons.
Note that, because the ASPs comprising several types of convex polygons shown in the study of this appendix are based on Tile(1,1)11(1,1)( 1 , 1 ) shown in [9] and assumes that Tile(1,1)11(1,1)( 1 , 1 ) is the chiral aperiodic monotile shown in [10], it is not confirmed that these sets are actually ASPs. Thus, because [10] is still a preview version, it does not hold if Tile(1,1)11(1,1)( 1 , 1 ) is not a chiral aperiodic monotile.